OpenStudy (anonymous):

Please help! :o Explain how the graph of each given function is a transformation of the graph of y=x^2. A. y=x^2-5 B. y=-2x^2 C. y=(x-5)^2 D. y=(-3x)^2

OpenStudy (jdoe0001):

\(\bf y = A ( Bx + C )^n + D\) A = shrinks or expands the parent graph by A factor B = shrinks or expands the parent graph by \(B^n\) factor C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D < 0, down

OpenStudy (jdoe0001):

 and so on